# Spin mechanism of drag resistance in strongly-correlated electron liquids

### v1

@@ Acknowledgments
We are grateful to Anton Andreev, Steven Kivelson, and Boris Spivak for insightful discussions. This research project was financially supported by the National Science Foundation (NSF) Grant No. DMR-2203411 and H. I. Romnes Faculty Fellowship provided by the University of Wisconsin-Madison Office of the Vice Chancellor for Research and Graduate Education with funding from the Wisconsin Alumni Research Foundation. This paper was finalized at the Aspen Center for Physics, during the program “Quantum Matter Through the Lens of Moiré Materials”, which is supported by the NSF Grant No. PHY-2210452.

### v2

@@ Acknowledgments
We are grateful to Anton Andreev, Steven Kivelson, and Boris Spivak for insightful discussions. This research project was financially supported by the National Science Foundation (NSF) Grant No. DMR-2203411 and H. I. Romnes Faculty Fellowship provided by the University of Wisconsin-Madison Office of the Vice Chancellor for Research and Graduate Education with funding from the Wisconsin Alumni Research Foundation. I.E. was supported by the University of Wisconsin – Madison. This paper was finalized at the Aspen Center for Physics, during the program “Quantum Matter Through the Lens of Moiré Materials”, which is supported by the NSF Grant No. PHY-2210452.